# Extremal distributions of partially hyperbolic systems: the Lipschitz threshold

### 01 Apr 2026 v1

@@ Abstract.
This gain in regularity has several applications to rigidity problems. In particular, we study the relationship between the $\ell$-integrability condition introduced by Eskin–Potrie–Zhang [27] and joint integrability in the conservative setting, yielding rigidity results for $u$-Gibbs measures.
We also obtain a $C^{\infty}$ classification of $3$-dimensional conservative partially hyperbolic diffeomorphisms with Lipschitz distributions, thereby answering a question of Carrasco–Hertz–Pujals [21] in the conservative setting under minimal regularity assumptions.
@@ 1.1. A brief introduction to partially hyperbolic systems
Partially hyperbolic systems arise naturally in smooth dynamics as a robust generalization of uniformly hyperbolic (Anosov) systems, capturing many of their key features such as stability and rich ergodic behavior. Recall that a diffeomorphism $f\colon M\to M$ is called partially hyperbolic if the tangent bundle admits a $Df$-invariant splitting
@@ 1.2. Extremal distributions and the Foulon–Hasselblatt cocycle
It is well known (see, e.g., [40, 54]) that the bundles $E^{s}$ and $E^{u}$ (and hence $E^{s}\oplus E^{u}$) are always Hölder continuous. On the other hand, higher regularity such as Lipschitz or $C^{1}$ is expected to be exceptional and to impose strong restrictions on the dynamics. This phenomenon was established by Foulon–Hasselblatt [29] for $3$-dimensional conservative $C^{r}$ Anosov flows. They showed that once $E^{s}\oplus E^{u}$ is Lipschitz, then $E^{s}\oplus E^{u}$ is automatically $C^{r-1}$, and the flow is either contact or (constant roof) suspension. In higher dimensions, Ghys [33] showed that for a $C^{2}$ Anosov flow, $C^{1}$ regularity of $E^{s}\oplus E^{u}$ forces the existence of a global cross section with constant return time.
@@
    the center cocycle $x\mapsto Df(x)|_{E^{c}}$ is cohomologous to a constant $\lambda^{c}$. Besides, $\lambda^{c}\neq 1$ if and only if $f$ is an Anosov diffeomorphism for which $E^{s}\oplus E^{u}$ is integrable.
@@
    • an isometric extension of an Anosov diffeomorphism;
@@
Remark $1.$
The regularity assumption on $E^{s}\oplus E^{u}$ is essentially optimal; see Example 1.4. Moreover, one cannot bootstrap the regularity of $E^{c}$ when $E^{s}\oplus E^{u}$ is integrable; see Example 1.5.
We also note that Theorem A admits a finite-regularity counterpart: if $f$ is $C^{r}$ ($r\geq 5$) and $E^{s}\oplus E^{u}$ is Lipschitz, then $E^{s}\oplus E^{u}$ is in fact $C^{r-1}$; see Remark 4.13 for more details.
@@ 1.3. Applications to the rigidity of $u$-Gibbs states
A key difficulty in applying this criterion is to express the QNI condition in geometric terms. This was achieved by Eskin–Potrie–Zhang [27], who showed that the failure of $\ell$-integrability (see Definition 6.2) implies the QNI condition. In particular, given a partially hyperbolic $u$-Gibbs state $\mu$ such that $E^{s}\oplus E^{u}$ is not Lipschitz on its support, then the support of $\mu$ is not $\ell$-integrable for any $\ell>2$ (see Lemma 6.5) and hence $\mu$ satisfies QNI (see [27, Theorem 8.1]).
@@
Since $\ell$-integrability for $\ell>2$ implies that $E^{s}\oplus E^{u}$ is Lipschitz, in the conservative setting, Theorem A yields:
@@ 1.4. Applications to the classification of 3D partially hyperbolic diffeomorphisms
Under additional assumptions, several positive results have been obtained. When the invariant distributions are $C^{\infty}$ and the pointwise Lyapunov exponents are constant, Carrasco–Pujals–Hertz [21] obtain a $C^{\infty}$ classification. It was later observed by Allout–Moghaddamfar [1] that this classification remains valid when the invariant distributions are merely $C^{1}$. Mion-Mouton [51] also obtains a $C^{\infty}$ classification when the invariant distributions are $C^{\infty}$ and $E^{s}\oplus E^{u}$ carries a contact structure.
Our main result shows that a surprisingly weak regularity assumption already forces full smooth rigidity in the volume-preserving setting.
@@
  • an isometric extension of an Anosov diffeomorphism of $\mathbb{T}^{2}$;
@@
Note that by Theorem A, the Lipschitz regularity of $E^{s}\oplus E^{u}$ already implies dynamical coherence, so the additional assumption concerns only the absolute continuity of $\mathcal{W}^{c}$. Moreover, the above regularity assumptions are essentially optimal; see Examples 1.4 and 1.5.
@@ 1.5. Further remarks and examples
Point (3) in Theorem A is related to a result of Gan–Shi [30, Theorem 1.1]. They show that for a $C^{1+\alpha}$ ($\alpha>0$) volume-preserving partially hyperbolic diffeomorphism on $\mathbb{T}^{3}$ homotopic to an Anosov automorphism, if $E^{s}\oplus E^{u}$ is integrable (and hence Lipschitz; see e.g. [45]), then the cocycle $x\mapsto Df(x)|_{E^{c}}$ is cohomologous to a constant $\lambda^{c}$. Our result extends this phenomenon to a broader setting: assuming only that $E^{s}\oplus E^{u}$ is Lipschitz, without any integrability or topological assumptions, we obtain the same cohomological rigidity. This suggests that the rigidity of the center dynamics can already be detected at the level of the regularity of $E^{s}\oplus E^{u}$.
@@
In our setting, we also obtain smoothness of the center bundle $E^{c}$ (when $f$ is accessible in Theorem A), but through a different mechanism: we exploit the smoothness of $E^{s}\oplus E^{u}$ together with the associated contact structure (see Corollary 3.4), thereby yielding smoothness of $E^{c}$.
@@
Example 1.7.
Even if $E^{s}\oplus E^{u}$ is $C^{\infty}$, the individual bundles $E^{s}$ and $E^{u}$ need not be $C^{2}$. A simple example is given by the product map $f=g\times\mathrm{Id}_{\mathbb{T}^{1}}$, where $g$ is a nonlinear volume-preserving Anosov diffeomorphism on $\mathbb{T}^{2}$. This shows that the conclusion in (2) of Theorem A is sharp.
@@ Acknowledgments.
We are grateful to Christian Bonatti, Sylvain Crovisier, Danijela Damjanovic, Ziqiang Feng, Shaobo Gan, Andrey Gogolev, Carlos Matheus, Federico Rodriguez Hertz, Yi Shi, Masato Tsujii, Amie Wilkinson, Danyu Zhang, and Zhiyuan Zhang for many helpful discussions.
@@ 2.1. Contact form and Reeb vector field
If $E$ is co-orientable, i.e. there exists a continuous $1$-dimensional distribution $F$ transverse to $E$, then there exists a $C^{r}$ nowhere-vanishing $1$-form $\alpha$ [^2] The choice of $\alpha$ is not unique: $\varphi\alpha$ is also a 1-form satisfying the condition for any nowhere vanishing Lipschitz function $\varphi$. such that $\ker\alpha=E$.
@@
In particular, if $E$ is a Lipschitz distribution, the 1-form $\alpha$ is Lipschitz, and is differentiable almost everywhere. The differential $d\alpha$ is a well-defined measurable ($L^{\infty}$) 2-form. The Frobenius theorem about integrablity of distributions also holds for Lipschitz distributions[29, Theorem 3].
Proposition 2.2 (see [59, Corollary A.3]).
@@ 2.2. Oseledets theorem and Lyapunov exponents
The following well-known result allows one to deduce uniform growth of cocycles from knowledge about exponents for every invariant measure. The proof is a corollary of a classical result on subadditive sequences (see [58] or [43, Chapter 4]).
Lemma 2.6 (see e.g. [23, Lemma 10]).
@@ 2.3. Partially hyperbolic systems on 3-manifolds
Let $f\colon M\to M$ be a $C^{1}$ diffeomorphism of a closed 3-manifold $M$. We say that $f$ is partially hyperbolic if there exists a $Df$-invariant splitting of the tangent bundle $TM=E^{s}\oplus E^{c}\oplus E^{u}$ such that
@@
Proposition 2.7 (see [39, Theorem 3.4], [28, Corollary 1.1]).
@@
The lemma is proved by Kalinin-Sadovskaya in a more general form [44, Lemma 4.1] via a beautiful application of Journé’s regularity lemma [42]. ∎
@@ 2.4. Cohomological equation and Livshits theorem
We then present a Livshits type theorem for a partially hyperbolic diffeomorphism on a 3-manifold, which is a special case of [62, Theorem A].
@@
Let $f\colon M\to M$ be a volume-preserving partially hyperbolic diffeomorphism on a 3-manifold. Let $U\subset M$ be an open accessibility class such that $f|_{U}$ is bijective and ergodic.
@@
    If $f$ is $C^{2}$ and $\phi$ is Hölder continuous, and there exists a measurable solution $\Phi$ of equation (1), then there exists a continuous solution $\tilde{\Phi}$ such that $\tilde{\Phi}=\Phi$ almost everywhere.
@@
The original statement [62, Theorem A] is given for the case where $U=M$. As noted in [62, Section 12], the proof applies to partially hyperbolic diffeomorphisms on non-compact manifolds provided that the inequalities in the definitions of partial hyperbolicity and strong $r$-bunching hold uniformly. Therefore, the theorem extends directly to $f$-invariant open accessibility classes as above.
@@ 3. Lipschitz regularity of $E^{s}\oplus E^{u}$ and center dynamics
Therefore, either $A^{+}$ or $A^{-}$ have full measure. Without loss of generality, we assume that $A^{+}$ has full measure. Then by Lemma 3.1, $\sqrt{|h(x)|}$ is a measurable solution to the cohomological equation (3). By Proposition 2.13, there is a continuous function $\tilde{h}\colon U\to{\mathbb{R}}^{+}$ solving (3), with $\tilde{h}=|h|$ for $m$-almost every $x\in U$. ∎

### 16 Apr 2026 v2

@@ Abstract.
This gain in regularity has several applications to rigidity problems. In particular, we study the relationship between the $\ell$-integrability condition introduced by Eskin–Potrie–Zhang [28] and joint integrability in the conservative setting, yielding rigidity results for $u$-Gibbs measures. We also obtain several $C^{\infty}$ classification results for partially hyperbolic diffeomorphisms on $3$-manifolds under various assumptions.
@@ 1.1. A brief introduction to partially hyperbolic systems
Partially hyperbolic systems arise naturally in smooth dynamics as a robust generalization of uniformly hyperbolic (Anosov) systems, capturing many of their key features such as stability and rich ergodic behavior. Recall that a diffeomorphism $f\colon M\to M$ is called partially hyperbolic if the tangent bundle admits a continuous $Df$-invariant splitting
@@ 1.2. Extremal distributions and the Foulon–Hasselblatt cocycle
It is well known (see, e.g., [42, 56]) that the bundles $E^{s}$ and $E^{u}$ (and hence $E^{s}\oplus E^{u}$) are always Hölder continuous. On the other hand, higher regularity such as Lipschitz or $C^{1}$ is expected to be exceptional and to impose strong restrictions on the dynamics. This phenomenon was established by Foulon–Hasselblatt [30] for $3$-dimensional conservative $C^{r}$ Anosov flows. They showed that once $E^{s}\oplus E^{u}$ is Lipschitz, then $E^{s}\oplus E^{u}$ is automatically $C^{r-1}$, and the flow is either contact or (constant roof) suspension. In higher dimensions, Ghys [35] showed that for a codimension one $C^{2}$ Anosov flow, $C^{1}$ regularity of $E^{s}\oplus E^{u}$ forces the existence of a global cross section with constant return time.
@@
    the center cocycle $x\mapsto Df(x)|_{E^{c}}$ is cohomologous to a constant $\lambda^{c}$. Besides, $\lambda^{c}\neq\pm 1$ if and only if $f$ is an Anosov diffeomorphism for which $E^{s}\oplus E^{u}$ is integrable.
@@
    • an isometric extension of a volume-preserving Anosov diffeomorphism on ${\mathbb{T}}^{2}$ with total space $M$ a Heisenberg nilmanifold;
@@
Remark 1.1.
Theorem A is sharp in the following sense:
  • The regularity assumption on $E^{s}\oplus E^{u}$ is essentially optimal; see Example 1.5.
  • The foliation $\mathcal{W}^{c}$ may fail to be absolutely continuous when $E^{s}\oplus E^{u}$ is integrable; see Example 1.6.
  • The distributions $E^{s}$ and $E^{u}$ may fail to be $C^{1+\mathrm{Lip}}$ even if $E^{s}\oplus E^{u}$ is $C^{\infty}$; see Example 1.7. In fact, if $E^{s}$ and $E^{u}$ are $C^{1+\mathrm{Lip}}$, then a stronger classification theorem can be obtained; see Corollary E.
Moreover, Theorem A admits a finite-regularity counterpart: if $f$ is $C^{r}$ ($r\geq 5$) and $E^{s}\oplus E^{u}$ is Lipschitz, then $E^{s}\oplus E^{u}$ is $C^{r-1}$; see Remark 4.13.
@@ 1.3. Applications to the rigidity of $u$-Gibbs states
A key difficulty in applying this criterion is to express the QNI condition in geometric terms. This was achieved by Eskin–Potrie–Zhang [28], who showed that the failure of $\ell$-integrability (see Definition 6.2) implies the QNI condition. In particular, given a partially hyperbolic $u$-Gibbs state $\mu$ such that $E^{s}\oplus E^{u}$ is not $C^{1}$ on its support, then the support of $\mu$ is not $\ell$-integrable for any $\ell\geq 2$ (see Lemma 6.5) and hence $\mu$ satisfies QNI (see [28, Theorem 8.1]).
@@
Since $\ell$-integrability for $\ell\geq 2$ implies that $E^{s}\oplus E^{u}$ is $C^{1}$. In the conservative setting, Theorem A yields:
@@ 1.4. Applications to the classification of 3D partially hyperbolic diffeomorphisms
Under additional assumptions, several positive results have been obtained. When the invariant distributions are $C^{\infty}$ and the pointwise Lyapunov exponents are constant, Carrasco–Pujals–Hertz [22] obtain a $C^{\infty}$ classification. It was later observed by Allout–Moghaddamfar [1] that this classification remains valid in the $C^{1}$ setting. Mion-Mouton [53] also obtains a $C^{\infty}$ classification when the invariant distributions are $C^{\infty}$ and $E^{s}\oplus E^{u}$ carries a contact structure.
By combining Theorem A with results of Avila–Viana–Wilkinson [7, 8], we obtain the following result, showing that in the volume-preserving setting, even a surprisingly weak regularity assumption is enough to force full smooth rigidity.
@@
  2. (2)
    an isometric extension of a volume-preserving Anosov diffeomorphism on ${\mathbb{T}}^{2}$;
  3. (3)
@@
Note that by Theorem A, the Lipschitz regularity of $E^{s}\oplus E^{u}$ already implies dynamical coherence, so the additional assumption concerns only the absolute continuity of $\mathcal{W}^{c}$. Moreover, the above regularity assumptions are essentially optimal; see Examples 1.5 and 1.6.
If one further assumes that $E^{s}$ and $E^{u}$ are $C^{1+\mathrm{Lip}}$, i.e., $C^{1}$ with Lipschitz continuous derivatives, then a stronger classification can be obtained.
Corollary E.
Let $f\colon M\to M$ be a $C^{\infty}$ volume-preserving partially hyperbolic diffeomorphism on a closed $3$-manifold. Assume that $E^{s}$ and $E^{u}$ are $C^{1+\mathrm{Lip}}$ and $\mathcal{W}^{c}$ is absolutely continuous; then $E^{s}$, $E^{u}$, and $E^{c}$ are all $C^{\infty}$. Up to a finite cover and an iterate $f$ is $C^{\infty}$-conjugate to one of the following:
  1. (1)
    an affine automorphism of a nilmanifold;
  2. (2)
    a constant roof suspension of a hyperbolic automorphism of ${\mathbb{T}}^{2}$;
  3. (3)
    the time-one map of a smooth time-change of the geodesic flow on a closed hyperbolic surface.
Remark 1.4.
The assumption that $E^{s}$ and $E^{u}$ are $C^{1+\mathrm{Lip}}$ is essentially optimal; see Remark 6.13. Moreover, the time-change in Case (3) is not arbitrary: it is a special algebraic time-change coming from a morphism from a lattice to the diagonal subgroup; see [34] for the precise description.
We also have a classification for contact partially hyperbolic diffeomorphisms on $3$-manifolds. Recall that a $C^{1}$ 11-form $\alpha$ is called contact if $\alpha\wedge d\alpha$ is non-vanishing everywhere.
Corollary F.
Let $f\colon M\to M$ be a $C^{\infty}$ partially hyperbolic diffeomorphism on a closed $3$-manifold preserving a $C^{1}$ contact form $\alpha$: $f^{*}\alpha=\alpha$. Then $\alpha$ is $C^{\infty}$, $f$ is volume-preserving, and the distributions $E^{s}\oplus E^{u}$ and $E^{c}$ are $C^{\infty}$. Up to a finite cover and an iterate, $f$ is $C^{\infty}$-conjugate to one of the following:
  • an isometric extension of a volume-preserving Anosov diffeomorphism on ${\mathbb{T}}^{2}$ with total space $M$ a Heisenberg nilmanifold;
  • the time-one map of a contact Anosov flow.
@@ 1.5. Further remarks and examples
Point (3) in Theorem A is related to a result of Gan–Shi [32, Theorem 1.1]. They show that for a $C^{1+\alpha}$ ($\alpha>0$) volume-preserving partially hyperbolic diffeomorphism on $\mathbb{T}^{3}$ homotopic to an Anosov automorphism, if $E^{s}\oplus E^{u}$ is integrable (and hence Lipschitz; see e.g. [47]), then the cocycle $x\mapsto Df(x)|_{E^{c}}$ is cohomologous to a constant $\lambda^{c}$. Our result extends this phenomenon to a broader setting: assuming only that $E^{s}\oplus E^{u}$ is Lipschitz, without any integrability or topological assumptions, we obtain the same cohomological rigidity. This suggests that the rigidity of the center dynamics can already be detected at the level of the regularity of $E^{s}\oplus E^{u}$.
@@
In our setting, we also obtain smoothness of the center bundle $E^{c}$ (when $f$ is accessible in Theorem A), but via a different mechanism: we exploit the smoothness of $E^{s}\oplus E^{u}$ together with the associated contact structure (see Corollary 3.5), which in turn yields the smoothness of $E^{c}$.
It is then natural to ask whether a similar phenomenon can occur when $E^{s}\oplus E^{u}$ is integrable and the individual distributions have higher regularity. Note that the assumption that $E^{s}\oplus E^{u}$ is $C^{\infty}$ alone is not sufficient; see Example 1.7.
Question 2.
Let $f\colon M\to M$ be a volume-preserving partially hyperbolic diffeomorphism on a $3$-manifold $M$. If $E^{s}\oplus E^{u}$ is integrable and both $E^{s}$ and $E^{u}$ are $C^{2}$, is $E^{c}$ absolutely continuous?
@@
Even if $E^{s}\oplus E^{u}$ is $C^{\infty}$, the individual bundles $E^{s}$ and $E^{u}$ need not be $C^{1+\mathrm{Lip}}$. One example is given by the product map $f=g\times\mathrm{Id}_{\mathbb{T}^{1}}$, where $g$ is a nonlinear volume-preserving Anosov diffeomorphism on $\mathbb{T}^{2}$. Other examples are provided by the non-algebraic contact Anosov flows constructed by Foulon–Hasselblatt [31], as well as Example 6.12 in this paper, for which $E^{s}$ and $E^{u}$ fail to be $C^{1+\mathrm{Lip}}$ by Corollary E.
Example 1.8.
@@ Acknowledgments.
We are grateful to Rafael Potrie for a very careful reading of an earlier version of this manuscript and for numerous helpful comments and suggestions that improved the paper. We are grateful to Amie Wilkinson for many helpful discussions and for posing a question that motivated Example 6.12. We are also grateful to Christian Bonatti, Sylvain Crovisier, Danijela Damjanovic, Ziqiang Feng, Shaobo Gan, Andrey Gogolev, Carlos Matheus, Federico Rodriguez Hertz, Yi Shi, Masato Tsujii, Danyu Zhang, and Zhiyuan Zhang for many helpful discussions.
@@ 2.1. Contact form and Reeb vector field
If $E$ is co-orientable, i.e. there exists a continuous $1$-dimensional distribution $F$ transverse to $E$, then there exists a $C^{r}$ nowhere-vanishing $1$-form $\alpha$ [^2] The choice of $\alpha$ is not unique: $\varphi\alpha$ is also a 1-form satisfying the condition for any nowhere vanishing $C^{r}$ function $\varphi$. such that $\ker\alpha=E$.
@@
In particular, if $E$ is a Lipschitz distribution, the 1-form $\alpha$ is Lipschitz, and is differentiable almost everywhere. The differential $d\alpha$ is a well-defined measurable ($L^{\infty}$) 2-form. The Frobenius theorem about integrablity of distributions also holds for Lipschitz distributions[30, Theorem 3].
Proposition 2.2 (see [61, Corollary A.3]).
@@ 2.2. Oseledets theorem and Lyapunov exponents
The following well-known result allows one to deduce uniform growth of cocycles from knowledge about exponents for every invariant measure. The proof is a corollary of a classical result on subadditive sequences (see [60] or [45, Chapter 4]).
Lemma 2.6 (see e.g. [24, Lemma 10]).
@@ 2.3. Partially hyperbolic systems on 3-manifolds
Let $f\colon M\to M$ be a $C^{1}$ diffeomorphism of a closed 3-manifold $M$. We say that $f$ is partially hyperbolic if there exists a continuous $Df$-invariant splitting of the tangent bundle $TM=E^{s}\oplus E^{c}\oplus E^{u}$ such that
@@
Proposition 2.7 (see [41, Theorem 3.4], [29, Corollary 1.1]).
@@
The lemma is proved by Kalinin-Sadovskaya in a more general form [46, Lemma 4.1] via a beautiful application of Journé’s regularity lemma [44]. ∎
@@ 2.4. Cohomological equation and Livshits theorem
We then present a Livshits type theorem for a partially hyperbolic diffeomorphism on a 3-manifold, which is a special case of [64, Theorem A].
@@
Let $f\colon M\to M$ be a partially hyperbolic diffeomorphism on a 3-manifold. Let $U\subset M$ be an open accessibility class such that $f|_{U}$ is bijective and ergodic.
@@
    If $f$ is $C^{2}$ and volume-preserving, $\phi$ is Hölder continuous, and there exists a measurable solution $\Phi$ of equation (1), then there exists a continuous solution $\tilde{\Phi}$ such that $\tilde{\Phi}=\Phi$ almost everywhere.
@@
The original statement [64, Theorem A] is given for the case where $U=M$. As noted in [64, Section 12], the proof applies to partially hyperbolic diffeomorphisms on non-compact manifolds provided that the inequalities in the definitions of partial hyperbolicity and strong $r$-bunching hold uniformly. Therefore, the theorem extends directly to $f$-invariant open accessibility classes as above.
@@ 3. Lipschitz regularity of $E^{s}\oplus E^{u}$ and center dynamics
Therefore, either $A^{+}$ or $A^{-}$ have full measure. Without loss of generality, we assume that $A^{+}$ has full measure. Then by Lemma 3.1, $\sqrt{h(x)}$ is a measurable solution to the cohomological equation (3). By Proposition 2.13, there is a continuous function $\tilde{h}\colon U\to{\mathbb{R}}^{+}$ solving (3), with $\tilde{h}=h$ for $m$-almost every $x\in U$. ∎
@@
Remark 3.4.
We will show later that if $f$ is as in Lemma 3.2, then $U=M$ (see Proposition 5.1). In particular, by Corollary 3.3, the center Lyapunov exponent of $f$ vanishes for every ergodic measure.